OpenStudy (anonymous):

what is the integral of 7y^2?

OpenStudy (anonymous):

\[7y^3/3 + C\] C is a constant

OpenStudy (anonymous):

\[\frac{7}{3}y^3\]

OpenStudy (anonymous):

7/3 y^3 +c?

OpenStudy (anonymous):

4!

OpenStudy (anonymous):

power rule: \[\int\limits x^n dx=\frac{x^{n+1}}{n+1}+c\] so \[\frac{7y^3}{3}+c\]

OpenStudy (anonymous):

\[\int\limits y ^{2}\] is y3/3 + c